OpenStudy (anonymous):

Integral of x(2x+3)^99 dx

zepdrix (zepdrix):

Expand the brackets first. Obviously that's the way to go. . . jk

OpenStudy (anonymous):

Substitution: let \(u=2x+3\), so that \(x=\dfrac{u-3}{2}\) and \(\dfrac{du}{2}=dx\). \[\large\int x(2x+3)^{99}~dx=\int\frac{u-3}{4}u^{99}~du=\int\left(\frac{u^{100}}{4}-\frac{3u^{99}}{4}\right)~du\]